# Simulation-based inference via telescoping ratio estimation for trawl processes

### v1

@@ 4.2 Neural point estimators
We begin by numerically deriving maximum likelihood estimates (MLE) from the TRE approximation. For comparison, we consider two benchmarks: MLEs obtained from standard NRE, and point estimates from the generalized method of moments (GMM). For GMM, the marginal parameters are inferred from the first four empirical moments and ACF parameters from the empirical autocorrelation function. Since these methods serve only as benchmarks, we provide their implementation details in Section LABEL:sm:point_estimators. Figure 3 illustrates the performance of these methods on a trawl process realization of length $1500$. Note that TRE achieves the closest match to both the true marginal distribution and the ACF structure.
@@
To confirm the superiority of TRE, we conduct a simulation study on $10^{4}$ samples $(\bm{x},\bm{\theta})$ drawn from the model. For each realization, we determine the point estimate via the BFGS optimization routine initialized at the true parameter $\bm{\theta}$, and summarize the estimation error in Table 1. We treat marginal and ACF parameters separately. For $(\mu,\sigma,\beta)$, we report the mean-absolute error (MAE), root mean-squared error (RMSE), and KL divergence between the true and inferred marginal distributions. For $\left(\gamma_{\textrm{acf}},\eta_{\textrm{acf}}\right)$, we directly compare the ACFs using the $L^{1}$ and $L^{2}$ distances rather than comparing individual parameters, as vastly different parameter values can sometimes yield nearly identical ACFs. While both TRE and NRE are trained on time series of length 1500, the LSTM encoder enables inference with time series of arbitrary length. Accross all metrics and lengths $k$, TRE substantially outperforms NRE and GMM. A natural question is how TRE compares to neural Bayes estimators (NBE), customized to target point summaries of the posterior distribution (Sainsbury-Dale et al. 2024). Training NBEs for trawl processes is tricky, as optimizing for the parameter-wise MAE or MSE fails due to the ACF weak-identifiability. We therefore train NBEs with multiple losses; see Section LABEL:sm:point_estimators for details. Table 1 shows the results for the NBE trained with $L^{2}$ ACF distance and MSE loss for marginal parameters. Even though NBEs have the advantage of directly optimizing the evaluation metric during training, we see that TRE achieves comparable or superior performance while providing the full posterior rather than just point estimates.
@@ 4.3 Calibration, coverage, and posterior checks
Next, Table 2 displays the BCE, $\mathcal{S}$, $\mathcal{B}$ and expected calibration error (ECE) metrics. We also report $W=\int_{0}^{1}\left|\mathcal{C}_{\alpha}-\alpha\right|\mathrm{d}\alpha$, which measures the overall deviation between the empirical and theoretical coverage. This metric corresponds to the Wasserstein distance between the distribution of posterior ranks and the uniform distribution, and can be interpreted as a rank check; see Section LABEL:SM:ranks. The calibrated TRE performs best across all metrics, with beta-calibration consistently improving individual classifier performance too. An interesting case is the NRE at $k=1000$, where the inherent underconfidence observed at $k=1500$ counterbalances, on average, the overconfidence gained from inputting shorter time series. However, the BCE and $\mathcal{S}$ metrics reveal that despite the reasonable coverage, the classifier’s quality is substantially worse than the corresponding TRE for $k=1000$.
@@
In this section, we employed beta-calibration to make a fair comparison between TRE and NRE. While isotonic regression is more flexible, it prevents gradient-based MCMC, thereby efficient posterior sampling and inference for the NRE. By contrast, the advocated Chebyshev polynomial approach is compatible with isotonic regression, and Section LABEL:SM:extended_beta_vs_iso shows that it can outperform beta-calibration for TRE. Methodologically, we note that the approximate densities $\theta^{i}\to\hat{p}(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1})$ may not integrate exactly to $1$. Sequential sampling in the dimensions of $\bm{\theta}$ may thus yield different results to those obtained by combining the approximate ratios within an MCMC, as the NRE components’ weights differ. We do not observe significant deviations at this stage, and refer this for future research.

### v2

@@ 4.2 Neural point estimators
We begin by numerically deriving maximum likelihood estimates (MLE) from the TRE approximation. For comparison, we consider two benchmarks: MLEs obtained from standard NRE, and point estimates from the generalized method of moments (GMM). For GMM, the marginal parameters are inferred from the first four empirical moments and ACF parameters from the empirical autocorrelation function. Since these methods serve only as benchmarks, we provide their implementation details in Section S5.1. Figure 3 illustrates the performance of these methods on a trawl process realization of length $1500$. Note that TRE achieves the closest match to both the true marginal distribution and the ACF structure.
@@
To confirm the superiority of TRE, we conduct a simulation study on $10^{4}$ samples $(\bm{x},\bm{\theta})$ drawn from the model. For each realization, we determine the point estimate via the BFGS optimization routine initialized at the true parameter $\bm{\theta}$, and summarize the estimation error in Table 1. We treat marginal and ACF parameters separately. For $(\mu,\sigma,\beta)$, we report the mean-absolute error (MAE), root mean-squared error (RMSE), and KL divergence between the true and inferred marginal distributions. For $\left(\gamma_{\textrm{acf}},\eta_{\textrm{acf}}\right)$, we directly compare the ACFs using the $L^{1}$ and $L^{2}$ distances rather than comparing individual parameters, as vastly different parameter values can sometimes yield nearly identical ACFs. While both TRE and NRE are trained on time series of length 1500, the LSTM encoder enables inference with time series of arbitrary length. Accross all metrics and lengths $k$, TRE substantially outperforms NRE and GMM. A natural question is how TRE compares to neural Bayes estimators (NBE), customized to target point summaries of the posterior distribution (Sainsbury-Dale et al. 2024). Training NBEs for trawl processes is tricky, as optimizing for the parameter-wise MAE or MSE fails due to the ACF weak-identifiability. We therefore train NBEs with multiple losses; see Section S5.1 for details. Table 1 shows the results for the NBE trained with $L^{2}$ ACF distance and MSE loss for marginal parameters. Even though NBEs have the advantage of directly optimizing the evaluation metric during training, we see that TRE achieves comparable or superior performance while providing the full posterior rather than just point estimates.
@@ 4.3 Calibration, coverage, and posterior checks
Next, Table 2 displays the BCE, $\mathcal{S}$, $\mathcal{B}$ and expected calibration error (ECE) metrics. We also report $W=\int_{0}^{1}\left|\mathcal{C}_{\alpha}-\alpha\right|\mathrm{d}\alpha$, which measures the overall deviation between the empirical and theoretical coverage. This metric corresponds to the Wasserstein distance between the distribution of posterior ranks and the uniform distribution, and can be interpreted as a rank check; see Section S4.2. The calibrated TRE performs best across all metrics, with beta-calibration consistently improving individual classifier performance too. An interesting case is the NRE at $k=1000$, where the inherent underconfidence observed at $k=1500$ counterbalances, on average, the overconfidence gained from inputting shorter time series. However, the BCE and $\mathcal{S}$ metrics reveal that despite the reasonable coverage, the classifier’s quality is substantially worse than the corresponding TRE for $k=1000$.
@@
In this section, we employed beta-calibration to make a fair comparison between TRE and NRE. While isotonic regression is more flexible, it prevents gradient-based MCMC, thereby efficient posterior sampling and inference for the NRE. By contrast, the advocated Chebyshev polynomial approach is compatible with isotonic regression, and Section S5.2 shows that it can outperform beta-calibration for TRE. Methodologically, we note that the approximate densities $\theta^{i}\to\hat{p}(\theta^{i}\mid\bm{x},\bm{\theta}^{1:i-1})$ may not integrate exactly to $1$. Sequential sampling in the dimensions of $\bm{\theta}$ may thus yield different results to those obtained by combining the approximate ratios within an MCMC, as the NRE components’ weights differ. We do not observe significant deviations at this stage, and refer this for future research.